6. What volume of 4.23 M Cs,Smot stuck solution is needed to make 280.0 mL of 0.781 M Cs 2 3 How many milliliter (mL of water needs to be added?
Answer #1

We have stock solution of 4.23 M Cs2S2O3 (M1) and we need to calculate the volume (V1)
required to make 280 mL (V2) of 0.781 M of Cs2S2O3 (M2);

So, we can use M1V1 = M2V2

V1 = M2V2/M1 = 280 (mL) x 0.781 (M)/ 4.23 (M) = 51.7 mL

We need 51.7 mL of Cs2S2O3 required to make 280 mL of 0.781 M Cs2S2O3 solution.
That means; the volume of water required = 280 - 51.7 = 228.3 mL

Therefore, we need to add 228.3 mL of water.